# Defending Against Prompt Injection With a Few DefensiveTokens

### v1

@@ 4.4. Analyzing the Utility-Security Trade-Off
Table 4. Utility (WinRate $\uparrow$) and security (ASR $\downarrow$) of test-time (TextGrad, Reminder, Sandwich, DefensiveToken) and training-time (StruQ, SecAlign using Full/LoRA fine-tuning) defense baselines. Numbers are averaged across models in Fig. 2 for visualization.
@@ 5. Conclusion
DefensiveToken effectively mitigates prompt injection while offering the system developer the flexibility to prioritize security or utility. Compared to other test-time defenses like Reminder or Sandwich defenses, DefensiveToken reduces attack success rate by two times to an order of magnitude. Compared to other defenses that require parameter fine-tuning like StruQ and SecAlign, DefensiveToken achieves a comparable level of robustness.
DefensiveToken only defends against prompt injections, where the user (instruction) is benign, and application-retrieved external data is malicious. DefensiveToken does not apply to other safety settings, e.g., preventing jailbreaks, system following attacks, and data extraction attacks, where the user is malicious.

### v2

@@ 4.5. Ablation Study
Multiple runs show that DefensiveTokens render reliable security. We do 5 runs when optimizing different 5 DefensiveTokens for Llama3.1-8B-Instruct, and study the randomness in the 9.1K-sized SEP benchmark. We get utility WinRate 53.84 ± 0.56 (53.49, 53.75, 54.43, 53.14, 54.38) and ASR 2.81 ± 1.09 (1.47, 4.22, 1.77, 3.79, 2.59).
@@ 5. Conclusion
DefensiveToken effectively mitigates prompt injection while offering the system developer the flexibility to prioritize security or utility. Compared to other test-time defenses like Reminder or Sandwich defenses, DefensiveToken reduces attack success rate by two times to an order of magnitude by asking the provider to optimize and release DefensiveTokens. Compared to other defenses that require parameter fine-tuning like StruQ and SecAlign, DefensiveToken achieves a comparable level of robustness.
DefensiveToken only defends against prompt injections, where the user (instruction) is benign, and application-retrieved external data is malicious. DefensiveToken does not apply to other safety settings, e.g., preventing jailbreaks, system following attacks, and data extraction attacks, where the user is malicious. Despite a reported trivial utility drop on AlpacaEval2 when applying DefensiveTokens, we do not know the utility on more labeled datasets, so our provided flexibility is still important, enabling developers to exclude DefensiveTokens when utility is prioritized.